NvidiaIntely (nvidiaintely):

A fossilized animal that was killed by a volcano eruption has 12.5 percent of its original amount of carbon-14 remaining; this radioisotope has a half-life of 5730 years. How long ago did the volcano erupt? 34,380 years ago 22,920 years ago 17,190 years ago 9,168 years ago

OpenStudy (kevin):

Nt = 12.5% No Nt/No = 12.5/100 Nt/No = (1/2)^(t / delta t) 12.5/100 = (1/2)^(t / 5730) t = 17190